1. what is the magnitude of the force on a charge of +2×10-7 C that is 0.3 from a charge of -5×10-7 C?
1
Expert's answer
2021-05-05T16:33:35-0400

"F=|\frac{kq_1q_2}{r^2}|,"

"F=\frac{9\cdot 10^9\cdot 2\cdot 10^{-7}\cdot 5\cdot 10^{-7}}{0.3^2\cdot 10^{-4}}=100~N."